OpenStudy (anonymous):

Solve the following equation: (-3e^2x)/(x^4) + (2e^2x)/(x^3) = 0 PLease help, I have no idea...

OpenStudy (anonymous):

\[\frac{ -3e ^{2x}}{ x^4 }+\frac{ 2e^{2x} }{x^3 }=0\] For your convenience.

OpenStudy (campbell_st):

well first you need a common denominator multiply the 2nd fraction by x and you get \[\frac{-3e^{2x}}{x^4} + \frac{2xe^{2x}}{x^4} = 0\] now factor the factor the numerator \[\frac{e^{2x}(2x -3)}{x^4} = 0\] so for the equation to equal zero, the the numerator must be zero so \[e^{2x}(2x - 3) = 0\] you need to know that the exponential will never be zero. \[e^{2x} \neq 0\] so the only way to get a zero numerator is to let 2x - 3 = 0 find the value of x, that makes it true...? hope it helps

OpenStudy (anonymous):

Thank you so much! I just forgot to take out the e^2x which made things really hard for me...